### v1

# criterion and density

Abstract.
For pure extensions $K=\mathbb{Q}(\alpha)$ with $\alpha^{n}=m$, we give a short proof, based only on Dedekind’s index theorem, of the $\alpha$-monogeneity criterion
$\mathbb{Z}[\alpha]=\mathcal{O}_{K}\iff\text{$m$ is square-free and }\nu_{p}(m^{p}-m)=1\ \text{for every prime }p\mid n.$
We then derive an explicit natural density $\delta_{n}=\frac{6}{\pi^{2}}\prod_{p\mid n}\frac{p}{p+1}$, independence across primes, refinements in arithmetic progressions, and discriminant-order asymptotics.
Key words and phrases:
monogenic fields; Dedekind’s index theorem; density in arithmetic progressions
2020 Mathematics Subject Classification:
11R04; 11R21; 11N37
MSC 2020. Primary 11R04; Secondary 11R21, 11N37.
Keywords. monogenic fields; Dedekind’s index theorem; density in arithmetic progressions.
1. Introduction
A basic task in algebraic number theory is to describe the ring of integers $\mathcal{O}_{K}$ of a number field $K$. We call $K$ monogenic if there exists an algebraic integer $\theta$ with $\mathcal{O}_{K}=\mathbb{Z}[\theta]$, so that every element of $\mathcal{O}_{K}$ can be expressed as a polynomial in a single generator, forming a power integral basis. This paper focuses the pure family $K=\mathbb{Q}(\alpha)$ with $\alpha^{n}=m$ and $m\in\mathbb{Z}$. We say that $K$ is $\alpha$-monogenic if $\mathbb{Z}[\alpha]=\mathcal{O}_{K}$. A natural question then arises: when is $K$ $\alpha$-monogenic? For example, in degree $2$ the answer depends on congruence conditions: for $K=\mathbb{Q}(\sqrt{a})$ one has
$\mathcal{O}_{K}=\mathbb{Z}[\sqrt{a}]\quad\text{if }a\equiv 2,3\pmod{4},\qquad\mathcal{O}_{K}=\mathbb{Z}\!\left[\frac{1+\sqrt{a}}{2}\right]\quad\text{if }a\equiv 1\pmod{4}$
(see, e.g., [Neu99, Ch.I]). Our paper gives an elementary criterion for a pure number field to be $\alpha$-monogenic, together with its natural density consequences.
1.1. Main results
Our first contribution is a one-page proof, relying only on Dedekind’s index theorem, which determines $\alpha$-monogeneity in every pure degrees.
Criterion for $\alpha$-monogeneity. Let $K=\mathbb{Q}(\alpha)$ with $\alpha^{n}=m$, $n\geq 2$, $m\in\mathbb{Z}\setminus\{0\}$. Then
$\mathcal{O}_{K}=\mathbb{Z}[\alpha]\ \Longleftrightarrow\ \text{$m$ is square-free and }\nu_{p}(m^{p}-m)=1\ \text{for every prime }p\mid n.$
The proof applies Dedekind’s index theorem to $f(X)=X^{n}-m$ and, writing $n=p^{r}s$ with $(p,s)=1$, compares $f$ with $g^{\,p^{r}}$ for $g(X)=X^{s}-m$ modulo $p$. A single $p$-adic evaluation in $(\mathbb{Z}/p^{2}\mathbb{Z})[X]/(g)$ shows that $p\mid[\mathcal{O}_{K}:\mathbb{Z}[\alpha]]$ iff $p^{2}\mid(m^{p^{r}}-m)$, which is equivalent to $\nu_{p}(m^{p}-m)\geq 2$, so the square-free condition controls primes dividing $m$. This yields, as immediate corollaries, the familiar residue tests in small degrees ($n=3,4,5,6$) without any integral-basis computation.
Our second contribution is a clean distribution theory for $\alpha$-monogenic pure fields. Let $\mathcal{E}_{p}=\{m\in\mathbb{Z}:\ m^{p}\equiv m\pmod{p^{2}}\}$. By Hensel, among units modulo $p^{2}$ the “bad” classes are exactly the (p1)(p-1) Teichmüller lifts. Conditioning on square-freeness, a proportion $1/(p+1)$ of $m$ lie in $\mathcal{E}_{p}$, and the events $\{m\in\mathcal{E}_{p}\}$ are independent across primes dividing $n$ (Chinese Remainder + uniform square-free distribution in classes). Consequently, for the set of $m$ with $x^{n}-m$ irreducible (a density-$1$ restriction by Eisenstein), the natural density of $\alpha$-monogeneity is
$\delta_{n}=\frac{6}{\pi^{2}}\prod_{p\mid n}\frac{p}{p+1},$
a product depending only on $\mathrm{rad}(n)$. We refine this density in arithmetic progressions (with explicit local factors), exhibit progressions on which $100\%$ of square-free $m$ are $\alpha$-monogenic, and count $\alpha$-monogenic pure fields by discriminant:
$N_{n}(Y)\sim\begin{cases}\delta_{n}\,(Y/n^{n})^{1/(n-1)},&n\ \text{odd},\\ 2\delta_{n}\,(Y/n^{n})^{1/(n-1)},&n\ \text{even}.\end{cases}$
1.2. Related results
There is an extensive literature on monogeneity, integral bases, and pure fields, among others, see Gaál’s monograph [Gaá19] for a comprehensive treatment and his recent survey [Gaá24] for an updated overview. More generally, for radical extensions $L(\sqrt[n]{\alpha})/L$, Smith proved a relative criterion for $\sqrt[n]{\alpha}$ to generate a power integral basis. When specialized to $L=\mathbb{Q}$ and the pure family $x^{n}-m$, this yields exactly the two local conditions in Theorem 2.6 (square-freeness of $m$ and $\nu_{p}(m^{p}-m)=1$ for all $p\mid n$) [Smi21]. Our proof has the advantage of being uniform and short in the pure case (one pp‑adic evaluation in $(\mathbb{Z}/p^{2}\mathbb{Z})[X]/(X^{s}-m)$). See also [Gas17, EF21] for earlier formulations in the case $x^{p^{r}}-a$ and the associated Wieferich‑type obstruction $a^{p}\equiv a\pmod{p^{2}}$.
A companion paper by the first-named author studies the full $p$-adic shape of explicit integral bases for pure fields and proves that, for each $p^{e}\parallel n$, the local shape is determined by $a\bmod p^{e+1}$, yielding a minimal global period $M(n)=\prod_{p^{e}\parallel n}p^{e+1}=n\cdot\mathrm{rad}(n)$. In that framework the same local invariant $r_{p}(a)=\nu_{p}(a^{p-1}-1)-1$ governs denominators and gives a Wieferich-type threshold at $p\mid n$, which is exactly the quantity appearing in our $\alpha$-monogeneity criterion $\nu_{p}(m^{p}-m)=1$. See [ND25] for details.
The density of square-free integers in progressions and basic $p$-adic features of Frobenius (Teichmüller lifts) will play a role in our distributional results, for accessible references see Brown [Bro21] and standard sources such as Cohen [Coh93] or Koblitz [Kob84]. For pure cubics, monogeneity (not necessarily via $\alpha$) and quantitative aspects were studied by Aygin–Nguyen [AN21]. In the monic coefficient space, Bhargava–Shankar–Wang [BSW22, BSW25] proved that a positive density $\zeta(2)^{-1}$ of polynomials (f) satisfy $\mathbb{Z}[x]/(f)=\mathcal{O}_{\mathbb{Q}[x]/(f)}$, and they determined the density of squarefree discriminants. They further extend these results to binary forms and resolve the arithmetic Bertini case of $\mathbb{P}^{1}_{\mathbb{Z}}$. Our work addresses instead the thin family $f_{m}(x)=x^{n}-m$: we give a complete local–global criterion for $\mathbb{Z}[\sqrt[n]{m}]$ to be maximal and compute the exact density $\frac{6}{\pi^{2}}\prod_{p\mid n}\frac{p}{p+1}$ along the m-line. Thus, BSW treat the generic coefficient space (via invariant‑theoretic methods), while we handle a rigid one‑parameter family (via Dedekind’s criterion), and the two sets of results are complementary. From a moduli‑theoretic perspective, schemes parameterizing monogenic generators and notions of local monogenicity are developed in [ABS23], while not used here, this situates our local conditions in a geometric framework.
1.3. Organization
Section 2 recalls preliminaries and gives the one-page proof of the criterion for $\alpha$-monogenic pure fields. Section 3 develops the density, independence, arithmetic-progression refinements, and discriminant counts.
1.4. Acknowledgement
The first author thanks the Morningside Center of Mathematics, Chinese Academy of Sciences, for its support and a stimulating research environment.
2. A criterion for $\alpha$-monogenic pure fields
To begin, we recall basic settings throughout the paper.
Definition 2.1.
An algebraic number field $K$ is called a pure number field if $K=\mathbb{Q}(\alpha)$, where $\alpha$ is a root of an irreducible polynomial $f(X)=X^{n}-m$ over $\mathbb{Q}$ with $m,n\in\mathbb{Z}$ and $n\geq 2$.
Definition 2.2.
If $\mathbb{Z}[\alpha]=\mathcal{O}_{K}$ for $K=\mathbb{Q}(\alpha)$, then $K$ is said to be $\alpha$-monogenic.
For example, when $K=\mathbb{Q}(\alpha)$ with $\alpha\in\{\sqrt{-2},\sqrt[3]{3},\sqrt[5]{6}\}$, one checks that $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$. We record a version of Dedekind’s index theorem as follows.
Theorem 2.3 (Dedekind’s index theorem).
Let $K=\mathbb{Q}(\alpha)$, where $\alpha$ is integral over $\mathbb{Z}$ with monic minimal polynomial $f\in\mathbb{Z}[X]$. For a prime $p$, write the factorization of $\overline{f}$ in $\mathbb{F}_{p}[X]$ as
$\overline{f}(X)=\overline{\pi}_{1}(X)^{e_{1}}\cdots\overline{\pi}_{g}(X)^{e_{g}},$
with pairwise distinct monic irreducibles $\overline{\pi}_{j}$. Let $\pi_{j}\in\mathbb{Z}[X]$ be monic lifts of $\overline{\pi}_{j}$ and define $F\in\mathbb{Z}[X]$ by
$f(X)=\pi_{1}(X)^{e_{1}}\cdots\pi_{g}(X)^{e_{g}}+p\,F(X).$
Then $p\mid(\mathcal{O}_{K}:\mathbb{Z}[\alpha])$ if and only if $\overline{\pi}_{j}\mid\overline{F}$ in $\mathbb{F}_{p}[X]$ for some $j$ with $e_{j}\geq 2$.
We also use the discriminant–index identity.
Theorem 2.4.
Let $K$ be a number field and let $\gamma_{1},\dots,\gamma_{m}$ be a $\mathbb{Z}$–basis of a sublattice $N\subset\mathcal{O}_{K}$ of finite index. Then
$D(\gamma_{1},\dots,\gamma_{m})=\bigl(\mathcal{O}_{K}:N\bigr)^{2}\,\mathrm{disc}(\mathcal{O}_{K}/\mathbb{Z}).$
In particular, if $K=\mathbb{Q}(\alpha)$ with $\alpha$ integral over $\mathbb{Z}$ and minimal polynomial $f$, then
(2.1) $\mathrm{disc}(f)=\bigl(\mathcal{O}_{K}:\mathbb{Z}[\alpha]\bigr)^{2}\,\mathrm{disc}(\mathcal{O}_{K}/\mathbb{Z}).$
Lemma 2.5.
Let $K=\mathbb{Q}(\alpha)$ with minimal polynomial $f(X)=X^{n}-m$. Then every prime dividing $\bigl(\mathcal{O}_{K}:\mathbb{Z}[\alpha]\bigr)$ divides $mn$. Equivalently, $K$ is $\alpha$-monogenic if and only if no prime $p\mid mn$ divides $\bigl(\mathcal{O}_{K}:\mathbb{Z}[\alpha]\bigr)$.
Proof.
Since $\mathrm{disc}(f)=(-1)^{\frac{n(n-1)}{2}}n^{n}m^{\,n-1}$, the claim follows from (2.1). ∎
We are now in a position to state and prove the main criterion of our work.
Theorem 2.6.
Let $K=\mathbb{Q}(\alpha)$ with $\alpha$ a root of the irreducible polynomial $f(X)=X^{n}-m\in\mathbb{Z}[X]$. Then $K$ is $\alpha$-monogenic, i.e. $\mathbb{Z}[\alpha]=\mathcal{O}_{K}$, if and only if
  • $m$ is square-free, and
  • $\nu_{p}(m^{p}-m)=1$ for every prime $p$ dividing $n$.
Proof.
By Lemma 2.5, only primes dividing $mn$ can divide the index.
Case $p\mid m$. Reducing $f$ modulo $p$ gives $\overline{f}(X)\equiv X^{n}$. Take $\pi(X)=X$ (multiplicity $e=n\geq 2$) and write
$f(X)=X^{n}+pF(X)\quad\text{with}\quad F(X)=-\frac{m}{p}\in\mathbb{Z}[X].$
By Theorem 2.3, we have $p\mid(\mathcal{O}_{K}:\mathbb{Z}[\alpha])$ if and only if $\overline{\pi}\mid\overline{F}$, i.e. iff $F\equiv 0\pmod{p}$, which holds exactly when $\nu_{p}(m)\geq 2$. Thus no such $p$ divides the index if and only if $m$ is square-free.
Case $p\mid n$. Write $n=p^{r}s$ with $(p,s)=1$ and set $g(X)=X^{s}-m$. Over $\mathbb{F}_{p}$ we have
$\overline{f}(X)=X^{n}-m\equiv\bigl(X^{s}-m\bigr)^{p^{r}}=\overline{g}(X)^{p^{r}},$
so every irreducible factor appears with multiplicity $\geq 2$. Define
$F(X)=\frac{f(X)-g(X)^{p^{r}}}{p}\in\mathbb{Z}[X],$
which is integral since $\binom{p^{r}}{k}\equiv 0\pmod{p}$ for $0<k<p^{r}$. Dedekind’s index theorem gives
$p\mid(\mathcal{O}_{K}:\mathbb{Z}[\alpha])\iff\overline{\pi}\mid\overline{F}$
for some irreducible factor $\overline{\pi}$ of $\overline{g}$. Let $\alpha_{0}$ be any root of $\overline{g}$ (hence of $\overline{f}$). In the ring $A=(\mathbb{Z}/p^{2}\mathbb{Z})[X]/(g)$ we have $X^{s}=m$, so
$\frac{f(X)-g(X)^{p^{r}}}{p}\equiv\frac{X^{n}-m}{p}\equiv\frac{m^{p^{r}}-m}{p}\pmod{pA}$.
Reducing modulo $p$ and evaluating at $\alpha_{0}$ yields
$\overline{F}(\alpha_{0})\equiv\frac{m^{p^{r}}-m}{p}\pmod{p}$.
Consequently, $p\mid(\mathcal{O}_{K}:\mathbb{Z}[\alpha])$ iff $p^{2}\mid(m^{p^{r}}-m)$, i.e. iff $\nu_{p}(m^{p^{r}}-m)\geq 2$, and hence
$p\nmid(\mathcal{O}_{K}:\mathbb{Z}[\alpha])\iff\nu_{p}(m^{p^{r}}-m)=1.$
Finally, if $p\mid m$ with $\nu_{p}(m)=1$ then $m^{p^{r}}-m=m(m^{p^{r}-1}-1)$ has $p$-adic valuation $1$. If $p\nmid m$, then
$\nu_{p}(m^{p^{r}}-m)=\nu_{p}\bigl(m(m^{p^{r}-1}-1)\bigr)=\nu_{p}(m^{p^{r}-1}-1)=\nu_{p}(m^{p-1}-1)=\nu_{p}(m^{p}-m),$
since $p\nmid(p^{r}-1)$ and the valuation is constant on exponents prime to $p$. Thus the condition is independent of $r$, and the two bullets in the statement are equivalent to $(\mathcal{O}_{K}:\mathbb{Z}[\alpha])=1$. ∎
Remark 2.7 (On the reduction from $p^{r}$ to $p$).
In the proof we used that, for any prime $p\mid n$ and $n=p^{r}s$ with $(p,s)=1$,
$\nu_{p}(m^{p^{r}}-m)=\nu_{p}(m^{p}-m).$
Indeed, if $p\nmid m$ then
$\nu_{p}(m^{p^{r}}-m)=\nu_{p}\!\big(m(m^{p^{r}-1}-1)\big)=\nu_{p}\!\big(m^{p^{r}-1}-1\big)=\nu_{p}\!\big(m^{p-1}-1\big)=\nu_{p}(m^{p}-m),$
since $p\nmid(p^{r}-1)$ and by the standard lifting-the-exponent (LTE) argument (see, e.g., [Coh93, Ch. 4] or [Kob84, §I.5]). If $p\mid m$ and $m$ is square-free at $p$, then $m^{p^{r}}-m=m\big(m^{p^{r}-1}-1\big)\equiv-m\not\equiv 0\pmod{p^{2}}$, whence $\nu_{p}(m^{p^{r}}-m)=1=\nu_{p}(m^{p}-m)$. This justifies the final reduction used in the proof.
We derive the following immediate consequences.
Corollary 2.8.
If $n=p^{r}$ with $p$ prime and $r\geq 1$, then $K$ is $\alpha$-monogenic if and only if $m$ is square-free and $\nu_{p}(m^{p}-m)=1$.
Corollary 2.9.
  • If $n=3$ then $K$ is $\alpha$-monogenic iff $m$ is square-free and $m\not\equiv\pm 1\pmod{9}$.
  • If $n=4$ then $K$ is $\alpha$-monogenic iff $m$ is square-free and $m\not\equiv 1\pmod{4}$.
  • If $n=5$ then $K$ is $\alpha$-monogenic iff $m$ is square-free and $m\not\equiv 1,7,18,24\pmod{25}$.
  • If $n=6$ then $K$ is $\alpha$-monogenic iff $m$ is square-free and $m\not\equiv 1\pmod{4}$ and $m\not\equiv\pm 1\pmod{9}$.
3. Natural density for $\alpha$-monogenic pure fields
In this section we compute the natural density of integers $m$ for which $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ when $\alpha^{n}=m$ is pure. Throughout, $n\geq 2$ is fixed.
Notation and preliminaries
For an integer $q\geq 1$ and $a\in\mathbb{Z}$, write
$\mathcal{A}(q,a)=\{m\in\mathbb{Z}:\ m\equiv a\pmod{q}\}$.
A set $E\subset\mathbb{Z}$ has natural density $\delta(E)$ if the limit
$\delta(E):=\lim_{X\to\infty}\frac{1}{X}\#\{m\in E:\ 1\leq m\leq X\}$
exists (equivalently, the two-sided limit with $|m|\leq X$ exists and has the same value for the congruence- and squarefreeness-defined sets we consider). We write $\mu$ for the Möbius function, and $\mu^{2}$ for the indicator of square-free integers.
We will also use the following elementary observations.
Lemma 3.1 (Square-free density in arithmetic progressions).
Let $q\geq 1$ and $a\in\mathbb{Z}$ with $\gcd(a,q)=1$. Then the set of square-free integers in the residue class $a\bmod q$ has natural density
$\delta\big(\{m:\ m\equiv a\bmod q,\ m\ \text{square-free}\}\big)=\frac{1}{q}\cdot\frac{1}{\zeta(2)}\cdot\prod_{\ell\mid q}\frac{1}{1-\ell^{-2}}.$
In particular, this density is the same for every $a$ coprime to $q$.
Proof.
Use $\mu^{2}(m)=\sum_{d^{2}\mid m}\mu(d)$ and interchange sums:
$\{m\leq X:\ m\equiv a\bmod q,\ \mu^{2}(m)=1\}=\sum_{\begin{subarray}{c}d\leq\sqrt{X}\\ \gcd(d,q)=1\end{subarray}}\mu(d)\ \#\{m\leq X:\ m\equiv a\bmod q,\ d^{2}\mid m\}.$
If $\gcd(d,q)>1$ and $\gcd(a,q)=1$, the inner count is zero. Otherwise, with $\gcd(d,q)=1$, the two conditions are equivalent to a single congruence modulo $qd^{2}$ by the Chinese Remainder Theorem, contributing $X/(qd^{2})+O(1)$. Hence
$\{m\leq X:\ m\equiv a\bmod q,\ \mu^{2}(m)=1\}=\frac{X}{q}\sum_{\gcd(d,q)=1}\frac{\mu(d)}{d^{2}}+O(\sqrt{X}).$
Since $\sum_{\gcd(d,q)=1}\mu(d)d^{-2}=\prod_{\ell\nmid q}(1-\ell^{-2})=\zeta(2)^{-1}\prod_{\ell\mid q}(1-\ell^{-2})^{-1}$, division by $X$ and passage to the limit yield the claim. ∎
Lemma 3.2 (Fixed points of Frobenius modulo $p^{2}$).
For a prime $p$ the congruence $x^{p}\equiv x\pmod{p^{2}}$ has exactly $p$ solutions modulo $p^{2}$: one is 0 and the others are the (p1)(p-1) Teichmüller lifts in $(\mathbb{Z}/p^{2}\mathbb{Z})^{\times}$. Moreover, each solution modulo $p$ lifts uniquely to a solution modulo $p^{k}$ for all $k\geq 1$.
Proof.
Let $g(x)=x^{p}-x$. In $\mathbb{F}_{p}$ the roots are all $x\in\mathbb{F}_{p}$, at each root mod $p$ one has $g^{\prime}(x)=px^{p-1}-1\equiv-1\pmod{p}$, which is invertible. By Hensel, each root lifts uniquely to $p^{k}$ for all $k$, so $\{x\bmod p^{2}:\ g(x)\equiv 0\}=p$. ∎
Remark 3.3 (Teichmüller lifts and $\mathcal{E}_{p}$).
The (p1)(p-1) unit solutions to $x^{p}\equiv x\pmod{p^{2}}$ are precisely the Teichmüller lifts of $\mathbb{F}_{p}^{\times}$. Thus $\mathcal{E}_{p}$ is the union of the single class $0\bmod p^{2}$ and these (p1)(p-1) unit classes (for $p=2$, interpret modulo $4$).
3.1. Local exclusions among square-free integers
For a fixed prime $p$, define the set of $p$-exceptional integers
$\mathcal{E}_{p}:=\{m\in\mathbb{Z}:\ m^{p}\equiv m\pmod{p^{2}}\}$.
By Lemma 3.2, $\mathcal{E}_{p}$ is the union of the single class $0\bmod p^{2}$ and the (p1)(p-1) unit-classes in $(\mathbb{Z}/p^{2}\mathbb{Z})^{\times}$ fixed by Frobenius. Intersecting with the square-free set automatically discards the class $0\bmod p^{2}$.
Proposition 3.4 (Local loss at a prime $p$).
Among square-free integers, the proportion that lie in $\mathcal{E}_{p}$ equals $\dfrac{1}{p+1}$. Equivalently, the surviving proportion at $p$ equals $\dfrac{p}{p+1}$.
Proof.
Among units modulo $p^{2}$ there are exactly (p1)(p-1) “bad” residue classes (the Teichmüller lifts). By Lemma 3.1, each such class contributes density
$\frac{1}{p^{2}}\cdot\frac{1}{\zeta(2)}\cdot\frac{1}{1-p^{-2}}=\frac{1}{\zeta(2)}\cdot\frac{1}{p^{2}-1}$
inside the square-free set. Summing over (p1)(p-1) classes, the lost fraction among square-free integers is
$\frac{p-1}{p^{2}-1}=\frac{1}{p+1},$
so the surviving fraction is $1-\frac{1}{p+1}=\frac{p}{p+1}$. ∎
3.2. Global density via independence across primes
Let $S$ denote the set of square-free integers. For a finite set of primes $P$ write
$\mathcal{B}_{P}\ :=\ S\cap\bigcap_{p\in P}\mathcal{E}_{p}.$
Note. The set $\bigcap_{p\in P}\mathcal{E}_{p}$ is a disjoint union of $\prod_{p\in P}(p-1)$ residue classes modulo $Q_{P}:=\prod_{p\in P}p^{2}$ (by the Chinese Remainder Theorem). Intersecting with $S$ is then handled by Lemma 3.1, applied class-by-class and summed.
Hence
(3.1) $\delta(\mathcal{B}_{P})=\frac{1}{\zeta(2)}\prod_{p\in P}\frac{p-1}{p^{2}-1}.$
Now inclusion–exclusion over $P\subseteq\{p:\ p\mid n\}$ shows that the set
$\mathcal{G}_{n}\ :=\ S\ \setminus\ \bigcup_{p\mid n}\mathcal{E}_{p}$
has density
$\delta(\mathcal{G}_{n})=\frac{1}{\zeta(2)}\prod_{p\mid n}\left(1-\frac{p-1}{p^{2}-1}\right)=\frac{1}{\zeta(2)}\prod_{p\mid n}\frac{p}{p+1}.$
This is the key combinatorial identity underlying the main density theorem below.
We also record that the restriction $x^{n}-m$ irreducible is asymptotically harmless as follows.
Lemma 3.5.
The set $\{m\in\mathbb{Z}:\ x^{n}-m\ \text{is irreducible over }\mathbb{Q}\}$ has natural density $1$.
Proof.
If $m$ has a prime factor $p$ with $\nu_{p}(m)=1$, then $x^{n}-m$ is Eisenstein at $p$, hence irreducible. Thus the complement is contained in $\{\pm 1\}$ and the set of squarefull integers $m=\pm\prod p_{i}^{e_{i}}$ with all $e_{i}\geq 2$. The number of squarefull integers $\leq X$ is $O(X^{1/2})$ (e.g. each squarefull $m$ can be written as $a^{2}b^{3}$, summing over $b\leq X^{1/3}$ gives $O(\sum_{b}X^{1/2}b^{-3/2})=O(X^{1/2})$). Hence the complement has density 0. ∎
We state and prove the second main theorem of the paper.
Theorem 3.6 (Natural density).
Fix $n\geq 2$. Among integers $m$ for which $x^{n}-m$ is irreducible, the set of $m$ with $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ has natural density
$\delta_{n}=\frac{6}{\pi^{2}}\prod_{p\mid n}\frac{p}{p+1}.$
In particular, $\delta_{n}$ depends only on $\mathrm{rad}(n)$.
Proof.
Recall that $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ iff $m$ is square-free and $\nu_{p}(m^{p}-m)=1$ for all $p\mid n$. Equivalently, the failure at $p$ is $m\in\mathcal{E}_{p}$. Therefore the desired set equals
$\{m:\ x^{n}-m\ \text{irreducible}\}\ \cap\ \mathcal{G}_{n}.$
By Lemma 3.5 the first set has density $1$, and by the computation preceding the theorem,
$\delta(\mathcal{G}_{n})=\frac{1}{\zeta(2)}\prod_{p\mid n}\frac{p}{p+1}=\frac{6}{\pi^{2}}\prod_{p\mid n}\frac{p}{p+1}.$
Intersecting with a density‑$1$ set does not change the density, which gives the claim. Since each local factor depends only on whether $p\mid n$, the product depends only on $\operatorname{rad}(n)$. ∎
Contrast with the coefficient‑box model. Over the full coefficient space, Bhargava–Shankar–Wang prove that a positive proportion $\zeta(2)^{-1}$ of monic polynomials (or binary nn‑ic forms) define maximal orders [BSW22, BSW25]. In our family $f_{m}(x)=x^{n}-m$, Theorem 3.6 gives the exact maximal‑order density $\delta_{n}=(6/\pi^{2})\prod_{p\mid n}\frac{p}{p+1}$.
Corollary 3.7 (Degree-specific criteria).
  1. (1)
    ($n=4$, pure quartics) $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ if and only if $m$ is square-free and $m\not\equiv 1\pmod{4}$.
  2. (2)
    ($n=6$, pure sextics) $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ if and only if $m$ is square-free, $m\not\equiv 1\pmod{4}$ and $m\not\equiv\pm 1\pmod{9}$.
Moreover, the corresponding densities are $\delta_{4}=\dfrac{4}{\pi^{2}}$ and $\delta_{6}=\dfrac{3}{\pi^{2}}$.
Proof.
Apply Theorem 2.6 with the primes dividing $n$.
(1) Here $p=2$. For any integer $m$, $\nu_{2}(m^{2}-m)=\nu_{2}\big(m(m-1)\big)$. If $m$ is even and square-free, then $\nu_{2}(m)=1$ and $\nu_{2}(m-1)=0$, hence $\nu_{2}(m^{2}-m)=1$. If $m$ is odd, then $\nu_{2}(m)=0$ and $\nu_{2}(m-1)=1$ iff $m\equiv 3\pmod{4}$. Thus among square-free $m$, the condition is $m\not\equiv 1\pmod{4}$. The density follows from Theorem 3.6 with $p=2$.
(2) Here $p\in\{2,3\}$. The $p=2$ condition is as in (1). For $p=3$ note $m^{3}-m=m(m-1)(m+1)$. If $3\mid m$ and $m$ is square-free, then $\nu_{3}(m^{3}-m)=1$. If $3\nmid m$, then $\nu_{3}(m^{3}-m)=1$ unless $m\equiv\pm 1\pmod{9}$, in which case $\nu_{3}\geq 2$. Combining gives the claimed congruences, and the density follows from Theorem 3.6. ∎
Remark 3.8 (Cubic case and field‑monogeneity).
For cube‑free $m$ and $p=3$, we have $\mathbb{Z}[\sqrt[3]{m}]=\mathcal{O}_{K}$ iff $m\not\equiv\pm 1\pmod{9}$. In the exceptional cases $m\equiv\pm 1\pmod{9}$, the field is nevertheless monogenic (though not via $\alpha$): one can take
$\theta=\frac{1\pm\alpha+\alpha^{2}}{3}.$
It is classical that $\theta$ is an algebraic integer in these residue classes and that $\mathcal{O}_{K}=\mathbb{Z}[\theta]$.
Related density in thin families. For a different thin one‑parameter family, Bardestani computes the density of primes $p$ for which $t^{q}-p$ is monogenic, with complementary phenomena in the cubic case [Bar14].
3.3. Consequences and applications of the density formula
Write $\delta_{n}=\frac{6}{\pi^{2}}\prod_{p\mid n}\frac{p}{p+1}$, we obtain the global counting with error term as follows.
Corollary 3.9.
Fix $n\geq 2$. Let
$\mathcal{M}_{n}(X)=\#\{\,|m|\leq X:\ x^{n}-m\text{ irreducible and }\mathcal{O}_{K}=\mathbb{Z}[\alpha]\,\}.$
Then
$\mathcal{M}_{n}(X)=2\delta_{n}\,X+O_{n}\!\left(X^{1/2}\right).$
If one restricts to $1\leq m\leq X$ the main term is $\delta_{n}X$ with the same error term.
Proof.
By Theorem 3.6 the relevant set has natural density $\delta_{n}$. The $O(X^{1/2})$ error follows from the proof of Lemma 3.1 (square-free counts in classes) combined with inclusion–exclusion over the finitely many primes $p\mid n$. ∎
The following result describes the arithmetic progressions together with their explicit local factors and depths.
Corollary 3.10.
Let $q\geq 1$ and $a\in\mathbb{Z}$ with $\gcd(a,q)=1$. Write $v_{p}(q)$ for the pp‑adic valuation. Then, among integers $m\equiv a\bmod q$ for which $x^{n}-m$ is irreducible, the set of $m$ with $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ has natural density
$\frac{1}{q}\cdot\frac{1}{\zeta(2)}\cdot\prod_{\ell\mid q}\frac{1}{1-\ell^{-2}}\ \times\ \prod_{p\mid n}\lambda_{p}(q,a),$
where the local factor at $p\mid n$ is
$\lambda_{p}(q,a)=\begin{cases}\displaystyle\frac{p}{p+1},&v_{p}(q)=0,\\[6.0pt] \displaystyle\frac{p-1}{p},&v_{p}(q)=1,\\[6.0pt] 1,&v_{p}(q)\geq 2\ \text{and}\ a\not\in\mathcal{E}_{p}\ \text{mod }p^{2},\\[4.0pt] 0,&v_{p}(q)\geq 2\ \text{and}\ a\in\mathcal{E}_{p}\ \text{mod }p^{2},\end{cases}$
with $\mathcal{E}_{p}=\{x\bmod p^{2}:\ x^{p}\equiv x\}$. For $p=2$ interpret this modulo $4$ (the unique bad unit class is $1\bmod 4$).
Proof.
By Lemma 3.1, among $m\equiv a\bmod q$ the square‑free density is $\frac{1}{q\zeta(2)}\prod_{\ell\mid q}(1-\ell^{-2})^{-1}$. Fix $p\mid n$. [^1] Uniformity among lifts from a fixed residue modulo $p$ to residues modulo $p^{2}$ follows by applying Lemma 3.1 to each class modulo $p^{2}$ and summing, using the Chinese Remainder Theorem to separate the $p$-part from the rest of the modulus.
If $v_{p}(q)=0$ there is no local information at $p$, so the survival probability is the global one $p/(p+1)$ (Proposition 3.4).
If $v_{p}(q)=1$, the congruence fixes $m\bmod p$ with $p\nmid a$. Among the $p$ lifts modulo $p^{2}$ above $a\bmod p$, exactly one is Teichmüller (bad), and square‑free integers are uniformly distributed across these lifts (by Lemma 3.1 with modulus $p^{2}$). Hence survival is (p1)/p(p-1)/p (Proposition 3.17).
If $v_{p}(q)\geq 2$, then $m\equiv a\bmod p^{2}$ is fixed. If $a\bmod p^{2}\in\mathcal{E}_{p}$, every such $m$ fails at $p$, otherwise every such $m$ survives at $p$. (For $p=2$ it suffices to work modulo $4$ for the valuation $\nu_{2}(m^{2}-m)$.)
Independence over $p$ (by the Chinese Remainder Theorem) yields the product of the local factors. ∎
Example 3.11 (Quartic case in a fixed progression).
Let $n=4$, $q=4$, $a\equiv 3\bmod 4$. Here $v_{2}(q)=2$ and $a\not\equiv 1\bmod 4$, so $\lambda_{2}(q,a)=1$. Therefore
$\delta\big(\{m\equiv 3\bmod 4:\ \mathcal{O}_{K}=\mathbb{Z}[\sqrt[4]{m}]\}\big)=\frac{1}{4}\cdot\frac{1}{\zeta(2)}\cdot\frac{1}{1-2^{-2}}=\frac{2}{\pi^{2}}.$
This reflects that on the class $3\bmod 4$ the $p=2$ constraint is automatically satisfied, only square‑freeness remains.
Remark 3.12 (Canonical modulus for refinement).
For questions beyond $\alpha$-monogeneity (e.g. tracking the full $p$-denominator profile of an integral basis), the canonical modulus is $M(n)=\prod_{p^{e}\parallel n}p^{e+1}=n\cdot\operatorname{rad}(n)$: the local $p$-shape depends only on $a\bmod p^{e+1}$ and this modulus is minimal. Our $\alpha$-monogeneity constraints at $p\mid n$ depend only on $a\bmod p^{2}$, but working modulo $M(n)$ aligns directly with the periodicity tables of the full shape, see [ND25].
The following result exhibits arithmetic progressions of square-free $m$ with $100\%$ $\alpha$-monogeneity.
Corollary 3.13.
Define
$M=\begin{cases}\displaystyle\prod_{\begin{subarray}{c}p\mid n\\ p\ \mathrm{odd}\end{subarray}}p^{2},&2\nmid n,\\[6.0pt] \displaystyle 4\prod_{\begin{subarray}{c}p\mid n\\ p\ \mathrm{odd}\end{subarray}}p^{2},&2\mid n.\end{cases}$
There exist residue classes $a\bmod M$ with $\gcd(a,M)=1$ such that every square-free $m\equiv a\bmod M$ satisfies $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$. Consequently, for each such $a$ the set $\{m\equiv a\bmod M:\ \mathcal{O}_{K}=\mathbb{Z}[\alpha]\}$ has density
$\frac{1}{M}\cdot\frac{1}{\zeta(2)}\cdot\prod_{\ell\mid M}\frac{1}{1-\ell^{-2}}.$
Proof.
Choose $a$ avoiding the finitely many bad classes modulo $M$: if $2\mid n$, avoid $1\bmod 4$ (and $\gcd(a,4)=1$ forces $a$ odd), and for each odd $p\mid n$ avoid the (p1)(p-1) Teichmüller classes modulo $p^{2}$. Then for any square-free $m\equiv a\bmod M$ one has $\nu_{p}(m^{p}-m)=1$ for all $p\mid n$. The density among $m\equiv a\bmod M$ is then given by Lemma 3.1. ∎
We obtain the following theorem, which gives the asymptotic count of $\alpha$-monogenic pure fields ordered by discriminant.
Theorem 3.14.
Let $N_{n}(Y)$ be the number of fields $K=\mathbb{Q}(\sqrt[n]{m})$ with $|\,\Delta_{K}\,|\leq Y$, $x^{n}-m$ irreducible, and $\mathcal{O}_{K}=\mathbb{Z}[\sqrt[n]{m}]$. Then, with $X=(Y/n^{n})^{1/(n-1)}$,
$N_{n}(Y)=\begin{cases}\ \delta_{n}\,X\ +\ O_{n}\!\big(X^{1/2}\big),&n\ \text{odd},\\[4.0pt] \ 2\delta_{n}\,X\ +\ O_{n}\!\big(X^{1/2}\big),&n\ \text{even}.\end{cases}$
Proof.
When $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$, the field discriminant equals the polynomial discriminant by (2.1): $\Delta_{K}=\mathrm{disc}(x^{n}-m)=\pm n^{n}m^{\,n-1}$. Thus $|\,\Delta_{K}\,|\leq Y$ is equivalent to $|m|\leq X$. For $n$ odd, the fields with parameters $m$ and m-m coincide (Indeed, when $n$ is odd, replacing $\alpha$ by $\alpha$ leaves $\mathbb{Q}(\alpha)$ unchanged, when $n$ is even, $\mathbb{Q}(\sqrt[n]{m})\neq\mathbb{Q}(\sqrt[n]{-m})$.) Hence counting positive $m$ gives distinct fields and yields the factor $\delta_{n}$, for $n$ even, $m\mapsto-m$ gives distinct fields, hence the factor $2\delta_{n}$. Apply Corollary 3.9. ∎
Immediate from $\delta_{n}=(6/\pi^{2})\prod_{p\mid n}p/(p+1)$ and the fact that adjoining a new prime factor multiplies by $p/(p+1)<1$ we obtain the following.
Example 3.15.
If $n\mid n^{\prime}$ then $\delta_{n^{\prime}}\leq\delta_{n}$. Explicitly,
$\delta_{2}=\delta_{4}=\frac{4}{\pi^{2}}\approx 0.4053,\qquad\delta_{3}=\frac{9}{2\pi^{2}}\approx 0.4559,\qquad\delta_{6}=\frac{3}{\pi^{2}}\approx 0.3040,\qquad\delta_{5}=\frac{5}{\pi^{2}}\approx 0.5066.$
The following result describes the failure distribution and independence phenomena.
Proposition 3.16.
Let $P=\{p:\ p\mid n\}$. Among square-free $m$, the events $\{m\in\mathcal{E}_{p}\}$ for $p\in P$ are independent. Hence, for any subset $S\subseteq P$,
[…]
$=\frac{6}{\pi^{2}}\Big(\prod_{p\in S}\frac{1}{p+1}\Big)\Big(\prod_{p\in P\setminus S}\frac{p}{p+1}\Big).$
In particular, the expected number of failing primes (among those dividing $n$) equals $\sum_{p\mid n}\frac{1}{p+1}$.
Proof.
By (3.1) with CRT, for any finite $P$ we have
$\delta\!\left(S\cap\bigcap_{p\in P}\mathcal{E}_{p}\right)=\frac{1}{\zeta(2)}\prod_{p\in P}\frac{p-1}{p^{2}-1}=\left(\frac{1}{\zeta(2)}\right)\prod_{p\in P}\left(\frac{1}{p+1}\cdot\frac{p}{p-1}\cdot\frac{p-1}{p}\right)=\delta(S)\prod_{p\in P}\frac{1}{p+1},$
so the events are independent within $S$ (and intersecting with the density‑1 irreducible set does not change the density). The stated formula and the expected value follow. ∎
Proposition 3.17 (Conditional local laws at a fixed prime).
Let $p\mid n$ and fix a residue class $a\!\!\pmod{p}$. Among square-free $m\equiv a\bmod p$, the survival probability at $p$ equals:
$\begin{cases}1,&a\equiv 0\ (\bmod p),\\[4.0pt] \frac{p-1}{p},&a\not\equiv 0\ (\bmod p).\end{cases}$
Consequently, conditioning on $m\equiv 0\pmod{p}$ neutralizes the $p$-constraint, while conditioning on any $a\not\equiv 0$ replaces the global factor $p/(p+1)$ by (p1)/p(p-1)/p.
Proof.
If $a\equiv 0$, then for square-free $m$ one has $v_{p}(m)=1$ and so $v_{p}(m^{p}-m)=1$ automatically. If $a\not\equiv 0$, then among the $p$ lifts modulo $p^{2}$ lying above $a$ exactly one is Teichmüller (bad), and by Lemma 3.1 each such lift carries the same square-free density, hence the survival fraction is (p1)/p(p-1)/p. ∎
Remark 3.18 (From $p$-level to $p^{2}$-level conditioning).
Proposition 3.17 conditions only on $m\bmod p$, giving survival probability (p1)/p(p-1)/p when $a\not\equiv 0\bmod p$. If one instead fixes $m\bmod p^{2}$, the survival probability becomes either $1$ (if $a\bmod p^{2}\notin\mathcal{E}_{p}$) or 0 (if $a\bmod p^{2}\in\mathcal{E}_{p}$). This is the mechanism behind Corollary 3.13: choosing a modulus $M$ divisible by $p^{2}$ for every $p\mid n$ (and $4$ for $p=2$) and selecting a good class modulo each $p^{2}$ forces the pp–constraint to hold identically on the progression.