### v2

# criterion and density

Abstract.
For pure extensions $K=\mathbb{Q}(\alpha)$ with $\alpha^{n}=m$, we give a short proof, based only on Dedekind’s index theorem, of the $\alpha$-monogeneity criterion
$\mathbb{Z}[\alpha]=\mathcal{O}_{K}\iff\text{$m$ is square-free and }\nu_{p}(m^{p}-m)=1\ \text{for every prime }p\mid n.$
We then derive an explicit natural density $\delta_{n}=\frac{6}{\pi^{2}}\prod_{p\mid n}\frac{p}{p+1}$, independence across primes, refinements in arithmetic progressions, and discriminant-order asymptotics.
Key words and phrases:
monogenic fields; Dedekind’s index theorem; density in arithmetic progressions
2020 Mathematics Subject Classification:
11R04; 11R21; 11N37
MSC 2020. Primary 11R04; Secondary 11R21, 11N37.
Keywords. monogenic fields; Dedekind’s index theorem; density in arithmetic progressions.
1. Introduction
A basic task in algebraic number theory is to describe the ring of integers $\mathcal{O}_{K}$ of a number field $K$. We call $K$ monogenic if there exists an algebraic integer $\theta$ with $\mathcal{O}_{K}=\mathbb{Z}[\theta]$, so that every element of $\mathcal{O}_{K}$ can be expressed as a polynomial in a single generator, forming a power integral basis. This paper focuses the pure family $K=\mathbb{Q}(\alpha)$ with $\alpha^{n}=m$ and $m\in\mathbb{Z}$. We say that $K$ is $\alpha$-monogenic if $\mathbb{Z}[\alpha]=\mathcal{O}_{K}$. A natural question then arises: when is $K$ $\alpha$-monogenic? For example, in degree $2$ the answer depends on congruence conditions: for $K=\mathbb{Q}(\sqrt{a})$ one has
$\mathcal{O}_{K}=\mathbb{Z}[\sqrt{a}]\quad\text{if }a\equiv 2,3\pmod{4},\qquad\mathcal{O}_{K}=\mathbb{Z}\!\left[\frac{1+\sqrt{a}}{2}\right]\quad\text{if }a\equiv 1\pmod{4}$
(see, e.g., [NEU99, Ch.I]). Our paper gives an elementary criterion for a pure number field to be $\alpha$-monogenic, together with its natural density consequences.
1.1. Main results
Our first contribution is a one-page proof, relying only on Dedekind’s index theorem, which determines $\alpha$-monogeneity in every pure degrees (see Theorem 2.7). More precisely, let $K=\mathbb{Q}(\alpha)$ with $\alpha^{n}=m$, $n\geq 2$, $m\in\mathbb{Z}\setminus\{0\}$. Then
$\mathcal{O}_{K}=\mathbb{Z}[\alpha]\ \Longleftrightarrow\ \text{$m$ is square-free and }\nu_{p}(m^{p}-m)=1\ \text{for every prime }p\mid n.$
The proof applies Dedekind’s index theorem to $f(X)=X^{n}-m$ and, writing $n=p^{r}s$ with $(p,s)=1$, compares $f$ with $g^{\,p^{r}}$ for $g(X)=X^{s}-m$ modulo $p$. A single $p$-adic evaluation in $(\mathbb{Z}/p^{2}\mathbb{Z})[X]/(g)$ shows that $p\mid[\mathcal{O}_{K}:\mathbb{Z}[\alpha]]$ iff $p^{2}\mid(m^{p^{r}}-m)$, which is equivalent to $\nu_{p}(m^{p}-m)\geq 2$, so the square-free condition controls primes dividing $m$. This yields, as immediate corollaries, the familiar residue tests in small degrees ($n=3,4,5,6$) without any integral-basis computation.
Our second contribution is a clean distribution theory for $\alpha$-monogenic pure fields (see Theorem 3.6). Let $\mathcal{E}_{p}=\{m\in\mathbb{Z}:\ m^{p}\equiv m\pmod{p^{2}}\}$. By Hensel, among units modulo $p^{2}$ the “bad” classes are exactly the (p1)(p-1) Teichmüller lifts. Conditioning on square-freeness, a proportion $1/(p+1)$ of $m$ lie in $\mathcal{E}_{p}$, and the events $\{m\in\mathcal{E}_{p}\}$ are independent across primes dividing $n$ (Chinese Remainder theorem and uniform square-free distribution in classes). Consequently, for the set of $m$ with $x^{n}-m$ irreducible, the natural density of $\alpha$-monogeneity is
$\delta_{n}=\frac{6}{\pi^{2}}\prod_{p\mid n}\frac{p}{p+1},$
a product depending only on $\mathrm{rad}(n)$. We refine this density in arithmetic progressions (with explicit local factors), exhibit progressions on which $100\%$ of square-free $m$ are $\alpha$-monogenic, and count $\alpha$-monogenic pure fields by discriminant (see Theorem 3.14)
$N_{n}(Y)\sim\begin{cases}\delta_{n}\,(Y/n^{n})^{1/(n-1)},&n\ \text{odd},\\ 2\delta_{n}\,(Y/n^{n})^{1/(n-1)},&n\ \text{even}.\end{cases}$
1.2. Related results
There is an extensive literature on monogeneity, integral bases, and pure fields. For comprehensive treatments, we refer to the monographs of Evertse–Győry [EG17] and Gaál [GAÁ19]. See also the recent surveys by Gaál [GAÁ24] and Evertse–Győry [EG24] for updated accounts of current developments.
More generally, for radical extensions $L(\sqrt[n]{\alpha})/L$, Smith in [SMI21] proved a relative criterion for $\sqrt[n]{\alpha}$ to generate a power integral basis. When specialized to $L=\mathbb{Q}$ and the pure family $x^{n}-m$, this yields exactly the two local conditions in Theorem 2.7 (square-freeness of $m$ and $\nu_{p}(m^{p}-m)=1$ for all $p\mid n$). Our proof has the advantage of being uniform and short in the pure case (one pp‑adic evaluation in $(\mathbb{Z}/p^{2}\mathbb{Z})[X]/(X^{s}-m)$). See also [GAS17, EL 21] for earlier formulations in the case $x^{p^{r}}-a$ and the associated Wieferich‑type obstruction $a^{p}\equiv a\pmod{p^{2}}$.
A companion paper  [NGU25] by the first-named author studies the full $p$-adic shape of explicit integral bases for pure fields and proves that, for each $p^{e}\parallel n$, the local shape is determined by $a\bmod p^{e+1}$, yielding a minimal global period $M(n)=\prod_{p^{e}\parallel n}p^{e+1}=n\cdot\mathrm{rad}(n)$. In that framework the same local invariant $r_{p}(a)=\nu_{p}(a^{p-1}-1)-1$ governs denominators and gives a Wieferich-type threshold at $p\mid n$, which is exactly the quantity appearing in our $\alpha$-monogeneity criterion $\nu_{p}(m^{p}-m)=1$.
The density of square-free integers in progressions and basic $p$-adic features of Frobenius (Teichmüller lifts) will play a role in our distributional results, for accessible references see Brown’s paper [BRO21] and standard sources such as Cohen’s book [COH93] or Koblitz’s book [KOB84]. For pure cubics, monogeneity and quantitative aspects were studied by Aygin–Nguyen [AN21]. In the monic coefficient space, Bhargava–Shankar–Wang [BSW22, BSW25] proved that a positive density $\zeta(2)^{-1}$ of polynomials $f$ satisfy $\mathbb{Z}[x]/(f)=\mathcal{O}_{\mathbb{Q}[x]/(f)}$, and they determined the density of squarefree discriminants. Our work addresses instead the thin family $f_{m}(x)=x^{n}-m$: we give a complete local–global criterion for $\mathbb{Z}[\sqrt[n]{m}]$ to be maximal and compute the exact density $\frac{6}{\pi^{2}}\prod_{p\mid n}\frac{p}{p+1}$. Thus, Bhargava–Shankar–Wang treat the generic coefficient space (via invariant‑theoretic methods), while we handle a rigid one‑parameter family (via Dedekind’s criterion), and the two sets of results are complementary. From a moduli‑theoretic perspective, schemes parameterizing monogenic generators and notions of local monogenicity are developed in [ABS23], while not used here, this situates our local conditions in a geometric framework.
1.3. Organization
Section 2 recalls preliminaries and gives the one-page proof of the criterion for $\alpha$-monogenic pure fields. Section 3 develops the density, independence, arithmetic-progression refinements, and discriminant counts.
1.4. Acknowledgement
The first author thanks the Morningside Center of Mathematics, Chinese Academy of Sciences, for its support and a stimulating research environment. We thank Professor Győry for his interest in our work and for drawing our attention to his monograph with Evertse, as well as to the updated version of [EG24]. We also thank Professor Gaál for several helpful communications.
2. A criterion for $\alpha$-monogenic pure fields
To begin, we recall basic settings throughout the paper.
Definition 2.1.
An algebraic number field $K$ is called a pure number field if $K=\mathbb{Q}(\alpha)$, where $\alpha$ is a root of an irreducible polynomial $f(X)=X^{n}-m$ over $\mathbb{Q}$ with $m,n\in\mathbb{Z}$ and $n\geq 2$.
Definition 2.2.
If $\mathbb{Z}[\alpha]=\mathcal{O}_{K}$ for $K=\mathbb{Q}(\alpha)$, then $K$ is said to be $\alpha$-monogenic.
For example, when $K=\mathbb{Q}(\alpha)$ with $\alpha\in\{\sqrt{-2},\sqrt[3]{3},\sqrt[5]{6}\}$, one checks that $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$.
Remark 2.3.
As Gaál kindly pointed out to us, the terminology $\alpha$-monogeneity is not standard in the literature. When a root $\alpha$ of a polynomial $f(x)$ generates a power integral basis, it is customary to say that the polynomial $f$ is monogenic, see, for example, the usage in Gassert’s work [GAS17]. We retain our terminology for clarity in the present context, but note this conventional viewpoint for the reader.
We record a version of Dedekind’s index theorem as follows.
Theorem 2.4 (Dedekind’s index theorem).
Let $K=\mathbb{Q}(\alpha)$, where $\alpha$ is integral over $\mathbb{Z}$ with monic minimal polynomial $f\in\mathbb{Z}[X]$. For a prime $p$, write the factorization of $\overline{f}$ in $\mathbb{F}_{p}[X]$ as
$\overline{f}(X)=\overline{\pi}_{1}(X)^{e_{1}}\cdots\overline{\pi}_{g}(X)^{e_{g}},$
with pairwise distinct monic irreducibles $\overline{\pi}_{j}$. Let $\pi_{j}\in\mathbb{Z}[X]$ be monic lifts of $\overline{\pi}_{j}$ and define $F\in\mathbb{Z}[X]$ by
$f(X)=\pi_{1}(X)^{e_{1}}\cdots\pi_{g}(X)^{e_{g}}+p\,F(X).$
Then $p\mid(\mathcal{O}_{K}:\mathbb{Z}[\alpha])$ if and only if $\overline{\pi}_{j}\mid\overline{F}$ in $\mathbb{F}_{p}[X]$ for some $j$ with $e_{j}\geq 2$.
We also use the discriminant–index identity.
Theorem 2.5.
Let $K$ be a number field and let $\gamma_{1},\dots,\gamma_{m}$ be a $\mathbb{Z}$–basis of a sublattice $N\subset\mathcal{O}_{K}$ of finite index. Then
$D(\gamma_{1},\dots,\gamma_{m})=\bigl(\mathcal{O}_{K}:N\bigr)^{2}\,\mathrm{disc}(\mathcal{O}_{K}/\mathbb{Z}).$
In particular, if $K=\mathbb{Q}(\alpha)$ with $\alpha$ integral over $\mathbb{Z}$ and minimal polynomial $f$, then
(2.1) $\mathrm{disc}(f)=\bigl(\mathcal{O}_{K}:\mathbb{Z}[\alpha]\bigr)^{2}\,\mathrm{disc}(\mathcal{O}_{K}/\mathbb{Z}).$
Lemma 2.6.
Let $K=\mathbb{Q}(\alpha)$ with minimal polynomial $f(X)=X^{n}-m$. Then every prime dividing $\bigl(\mathcal{O}_{K}:\mathbb{Z}[\alpha]\bigr)$ divides $mn$. Equivalently, $K$ is $\alpha$-monogenic if and only if no prime $p\mid mn$ divides $\bigl(\mathcal{O}_{K}:\mathbb{Z}[\alpha]\bigr)$.
Proof.
Since $\mathrm{disc}(f)=(-1)^{\frac{n(n-1)}{2}}n^{n}m^{\,n-1}$, the claim follows from (2.1). ∎
We are now in a position to state and prove the main criterion of our work.
Theorem 2.7.
Let $K=\mathbb{Q}(\alpha)$ with $\alpha$ a root of the irreducible polynomial $f(X)=X^{n}-m\in\mathbb{Z}[X]$. Then $K$ is $\alpha$-monogenic, i.e. $\mathbb{Z}[\alpha]=\mathcal{O}_{K}$, if and only if
  • $m$ is square-free, and
  • $\nu_{p}(m^{p}-m)=1$ for every prime $p$ dividing $n$.
Proof.
By Lemma 2.6, only primes dividing $mn$ can divide the index.
Case $p\mid m$. Reducing $f$ modulo $p$ gives $\overline{f}(X)\equiv X^{n}$. Take $\pi(X)=X$ (multiplicity $e=n\geq 2$) and write
$f(X)=X^{n}+pF(X)\quad\text{with}\quad F(X)=-\frac{m}{p}\in\mathbb{Z}[X].$
By Theorem 2.4, we have $p\mid(\mathcal{O}_{K}:\mathbb{Z}[\alpha])$ if and only if $\overline{\pi}\mid\overline{F}$, i.e. iff $F\equiv 0\pmod{p}$, which holds exactly when $\nu_{p}(m)\geq 2$. Thus no such $p$ divides the index if and only if $m$ is square-free.
Case $p\mid n$. Write $n=p^{r}s$ with $(p,s)=1$ and set $g(X)=X^{s}-m$. Over $\mathbb{F}_{p}$ we have
$\overline{f}(X)=X^{n}-m\equiv\bigl(X^{s}-m\bigr)^{p^{r}}=\overline{g}(X)^{p^{r}},$
so every irreducible factor appears with multiplicity $\geq 2$. Define
$F(X)=\frac{f(X)-g(X)^{p^{r}}}{p}\in\mathbb{Z}[X],$
which is integral since $\binom{p^{r}}{k}\equiv 0\pmod{p}$ for $0<k<p^{r}$. Dedekind’s index theorem gives
$p\mid(\mathcal{O}_{K}:\mathbb{Z}[\alpha])\iff\overline{\pi}\mid\overline{F}$
for some irreducible factor $\overline{\pi}$ of $\overline{g}$. Let $\alpha_{0}$ be any root of $\overline{g}$ (hence of $\overline{f}$). In the ring $A=(\mathbb{Z}/p^{2}\mathbb{Z})[X]/(g)$ we have $X^{s}=m$, so
$\frac{f(X)-g(X)^{p^{r}}}{p}\equiv\frac{X^{n}-m}{p}\equiv\frac{m^{p^{r}}-m}{p}\pmod{pA}$.
Reducing modulo $p$ and evaluating at $\alpha_{0}$ yields
$\overline{F}(\alpha_{0})\equiv\frac{m^{p^{r}}-m}{p}\pmod{p}$.
Consequently, $p\mid(\mathcal{O}_{K}:\mathbb{Z}[\alpha])$ iff $p^{2}\mid(m^{p^{r}}-m)$, i.e. iff $\nu_{p}(m^{p^{r}}-m)\geq 2$, and hence
$p\nmid(\mathcal{O}_{K}:\mathbb{Z}[\alpha])\iff\nu_{p}(m^{p^{r}}-m)=1.$
Finally, if $p\mid m$ with $\nu_{p}(m)=1$ then $m^{p^{r}}-m=m(m^{p^{r}-1}-1)$ has $p$-adic valuation $1$. If $p\nmid m$, then
$\nu_{p}(m^{p^{r}}-m)=\nu_{p}\bigl(m(m^{p^{r}-1}-1)\bigr)=\nu_{p}(m^{p^{r}-1}-1)=\nu_{p}(m^{p-1}-1)=\nu_{p}(m^{p}-m),$
since $p\nmid(p^{r}-1)$ and the valuation is constant on exponents prime to $p$. Thus the condition is independent of $r$, and the two bullets in the statement are equivalent to $(\mathcal{O}_{K}:\mathbb{Z}[\alpha])=1$. ∎
Remark 2.8 (On the reduction from $p^{r}$ to $p$).
In the proof we used that, for any prime $p\mid n$ and $n=p^{r}s$ with $(p,s)=1$,
$\nu_{p}(m^{p^{r}}-m)=\nu_{p}(m^{p}-m).$
Indeed, if $p\nmid m$ then
$\nu_{p}(m^{p^{r}}-m)=\nu_{p}\!\big(m(m^{p^{r}-1}-1)\big)=\nu_{p}\!\big(m^{p^{r}-1}-1\big)=\nu_{p}\!\big(m^{p-1}-1\big)=\nu_{p}(m^{p}-m),$
since $p\nmid(p^{r}-1)$ and by the standard lifting-the-exponent (LTE) argument (see, e.g., [COH93, Ch. 4] or [KOB84, §I.5]). If $p\mid m$ and $m$ is square-free at $p$, then $m^{p^{r}}-m=m\big(m^{p^{r}-1}-1\big)\equiv-m\not\equiv 0\pmod{p^{2}}$, whence $\nu_{p}(m^{p^{r}}-m)=1=\nu_{p}(m^{p}-m)$. This justifies the final reduction used in the proof.
We derive the following immediate consequences.
Corollary 2.9.
  • If $n=3$ then $K$ is $\alpha$-monogenic iff $m$ is square-free and $m\not\equiv\pm 1\pmod{9}$.
  • If $n=4$ then $K$ is $\alpha$-monogenic iff $m$ is square-free and $m\not\equiv 1\pmod{4}$.
  • If $n=5$ then $K$ is $\alpha$-monogenic iff $m$ is square-free and $m\not\equiv 1,7,18,24\pmod{25}$.
  • If $n=6$ then $K$ is $\alpha$-monogenic iff $m$ is square-free and $m\not\equiv 1\pmod{4}$ and $m\not\equiv\pm 1\pmod{9}$.
3. Natural density for $\alpha$-monogenic pure fields
In this section we compute the natural density of integers $m$ for which $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ when $K$ is pure. Throughout, $n\geq 2$ is fixed.
Notation and preliminaries
For an integer $q\geq 1$ and $a\in\mathbb{Z}$, write
$\mathcal{A}(q,a)=\{m\in\mathbb{Z}:\ m\equiv a\pmod{q}\}$.
A set $E\subset\mathbb{Z}$ has natural density $\delta(E)$ if the following limit exists
$\delta(E):=\lim_{X\to\infty}\frac{1}{X}\#\{m\in E:\ 1\leq m\leq X\}.$
We write $\mu$ for the Möbius function, and $\mu^{2}$ for the indicator of square-free integers. We will also use the following elementary observations.
Lemma 3.1 (Square-free density in arithmetic progressions).
Let $q\geq 1$ and $a\in\mathbb{Z}$ with $\gcd(a,q)=1$. Then the set of square-free integers in the residue class $a\bmod q$ has natural density
$\delta\big(\{m:\ m\equiv a\bmod q,\ m\ \text{square-free}\}\big)=\frac{1}{q}\cdot\frac{1}{\zeta(2)}\cdot\prod_{\ell\mid q}\frac{1}{1-\ell^{-2}}.$
In particular, this density is the same for every $a$ coprime to $q$.
Proof.
Use $\mu^{2}(m)=\sum_{d^{2}\mid m}\mu(d)$ and interchange sums:
$\{m\leq X:\ m\equiv a\bmod q,\ \mu^{2}(m)=1\}=\sum_{\begin{subarray}{c}d\leq\sqrt{X}\\ \gcd(d,q)=1\end{subarray}}\mu(d)\ \#\{m\leq X:\ m\equiv a\bmod q,\ d^{2}\mid m\}.$
If $\gcd(d,q)>1$ and $\gcd(a,q)=1$, the inner count is zero. Otherwise, with $\gcd(d,q)=1$, the two conditions are equivalent to a single congruence modulo $qd^{2}$ by the Chinese Remainder Theorem, contributing $X/(qd^{2})+O(1)$. Hence
$\{m\leq X:\ m\equiv a\bmod q,\ \mu^{2}(m)=1\}=\frac{X}{q}\sum_{\gcd(d,q)=1}\frac{\mu(d)}{d^{2}}+O(\sqrt{X}).$
Since $\sum_{\gcd(d,q)=1}\mu(d)d^{-2}=\prod_{\ell\nmid q}(1-\ell^{-2})=\zeta(2)^{-1}\prod_{\ell\mid q}(1-\ell^{-2})^{-1}$, division by $X$ and passage to the limit yield the claim. ∎
Lemma 3.2 (Fixed points of Frobenius modulo $p^{2}$).
For a prime $p$ the congruence $x^{p}\equiv x\pmod{p^{2}}$ has exactly $p$ solutions modulo $p^{2}$: one is 0 and the others are the (p1)(p-1) Teichmüller lifts in $(\mathbb{Z}/p^{2}\mathbb{Z})^{\times}$. Moreover, each solution modulo $p$ lifts uniquely to a solution modulo $p^{k}$ for all $k\geq 1$.
Proof.
Let $g(x)=x^{p}-x$. In $\mathbb{F}_{p}$ the roots are all $x\in\mathbb{F}_{p}$, at each root mod $p$ one has $g^{\prime}(x)=px^{p-1}-1\equiv-1\pmod{p}$, which is invertible. By Hensel’s lemma, each root lifts uniquely to $p^{k}$ for all $k$, so $\{x\bmod p^{2}:\ g(x)\equiv 0\}=p$. ∎
Remark 3.3 (Teichmüller lifts and $\mathcal{E}_{p}$).
The (p1)(p-1) unit solutions to $x^{p}\equiv x\pmod{p^{2}}$ are precisely the Teichmüller lifts of $\mathbb{F}_{p}^{\times}$. Thus $\mathcal{E}_{p}$ is the union of the single class $0\bmod p^{2}$ and these (p1)(p-1) unit classes (for $p=2$, interpret modulo $4$).
3.1. Local exclusions among square-free integers
For a fixed prime $p$, define the set of $p$-exceptional integers
$\mathcal{E}_{p}:=\{m\in\mathbb{Z}:\ m^{p}\equiv m\pmod{p^{2}}\}$.
By Lemma 3.2, $\mathcal{E}_{p}$ is the union of the single class $0\bmod p^{2}$ and the (p1)(p-1) unit-classes in $(\mathbb{Z}/p^{2}\mathbb{Z})^{\times}$ fixed by Frobenius. Intersecting with the square-free set automatically discards the class $0\bmod p^{2}$.
Proposition 3.4 (Local loss at a prime $p$).
Among square-free integers, the proportion that lie in $\mathcal{E}_{p}$ equals $\dfrac{1}{p+1}$. Equivalently, the surviving proportion at $p$ equals $\dfrac{p}{p+1}$.
Proof.
Among units modulo $p^{2}$ there are exactly (p1)(p-1) “bad” residue classes (the Teichmüller lifts). By Lemma 3.1, each such class contributes density
$\frac{1}{p^{2}}\cdot\frac{1}{\zeta(2)}\cdot\frac{1}{1-p^{-2}}=\frac{1}{\zeta(2)}\cdot\frac{1}{p^{2}-1}$
inside the square-free set. Summing over (p1)(p-1) classes, the lost fraction among square-free integers is
$\frac{p-1}{p^{2}-1}=\frac{1}{p+1},$
so the surviving fraction is $1-\frac{1}{p+1}=\frac{p}{p+1}$. ∎
3.2. Global density via independence across primes
Let $S$ denote the set of square-free integers. For a finite set of primes $P$ write
$\mathcal{B}_{P}\ :=\ S\cap\bigcap_{p\in P}\mathcal{E}_{p}.$
Note. The set $\bigcap_{p\in P}\mathcal{E}_{p}$ is a disjoint union of $\prod_{p\in P}(p-1)$ residue classes modulo $Q_{P}:=\prod_{p\in P}p^{2}$ (by the Chinese Remainder Theorem). Intersecting with $S$ is then handled by Lemma 3.1, applied class-by-class and summed.
Hence
(3.1) $\delta(\mathcal{B}_{P})=\frac{1}{\zeta(2)}\prod_{p\in P}\frac{p-1}{p^{2}-1}.$
Now inclusion–exclusion over $P\subseteq\{p:\ p\mid n\}$ shows that the set
$\mathcal{G}_{n}\ :=\ S\ \setminus\ \bigcup_{p\mid n}\mathcal{E}_{p}$
has density
$\delta(\mathcal{G}_{n})=\frac{1}{\zeta(2)}\prod_{p\mid n}\left(1-\frac{p-1}{p^{2}-1}\right)=\frac{1}{\zeta(2)}\prod_{p\mid n}\frac{p}{p+1}.$
This is the key combinatorial identity underlying the main density theorem below. We also record that the restriction $x^{n}-m$ irreducible is asymptotically harmless as follows.
Lemma 3.5.
The set $\{m\in\mathbb{Z}:\ x^{n}-m\ \text{is irreducible over }\mathbb{Q}\}$ has natural density $1$.
Proof.
If $m$ has a prime factor $p$ with $\nu_{p}(m)=1$, then $x^{n}-m$ is Eisenstein at $p$, hence irreducible. Thus the complement is contained in $\{\pm 1\}$ and the set of squarefull integers $m=\pm\prod p_{i}^{e_{i}}$ with all $e_{i}\geq 2$. The number of squarefull integers $\leq X$ is $O(X^{1/2})$ (e.g. each squarefull $m$ can be written as $a^{2}b^{3}$, summing over $b\leq X^{1/3}$ gives $O(\sum_{b}X^{1/2}b^{-3/2})=O(X^{1/2})$). Hence the complement has density 0. ∎
We state and prove the second main theorem of the paper.
Theorem 3.6 (Natural density).
Fix $n\geq 2$. Among integers $m$ for which $x^{n}-m$ is irreducible, the set of $m$ with $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ has natural density
$\delta_{n}=\frac{6}{\pi^{2}}\prod_{p\mid n}\frac{p}{p+1}.$
In particular, $\delta_{n}$ depends only on $\mathrm{rad}(n)$.
Proof.
Recall that $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ iff $m$ is square-free and $\nu_{p}(m^{p}-m)=1$ for all $p\mid n$. Equivalently, the failure at $p$ is $m\in\mathcal{E}_{p}$. Therefore the desired set equals
$\{m:\ x^{n}-m\ \text{irreducible}\}\ \cap\ \mathcal{G}_{n}.$
By Lemma 3.5 the first set has density $1$, and by the computation preceding the theorem,
$\delta(\mathcal{G}_{n})=\frac{1}{\zeta(2)}\prod_{p\mid n}\frac{p}{p+1}=\frac{6}{\pi^{2}}\prod_{p\mid n}\frac{p}{p+1}.$
Intersecting with a density‑$1$ set does not change the density, which gives the claim. Since each local factor depends only on whether $p\mid n$, the product depends only on $\operatorname{rad}(n)$. ∎
Below are direct consequences of the density theorem.
Corollary 3.7 (Degree-specific criteria).
  1. (1)
    ($n=4$, pure quartics) $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ if and only if $m$ is square-free and $m\not\equiv 1\pmod{4}$.
  2. (2)
    ($n=6$, pure sextics) $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ if and only if $m$ is square-free, $m\not\equiv 1\pmod{4}$ and $m\not\equiv\pm 1\pmod{9}$.
Moreover, the corresponding densities are $\delta_{4}=\dfrac{4}{\pi^{2}}$ and $\delta_{6}=\dfrac{3}{\pi^{2}}$.
Proof.
Apply Theorem 2.7 with the primes dividing $n$.
(1) Here $p=2$. For any integer $m$, $\nu_{2}(m^{2}-m)=\nu_{2}\big(m(m-1)\big)$. If $m$ is even and square-free, then $\nu_{2}(m)=1$ and $\nu_{2}(m-1)=0$, hence $\nu_{2}(m^{2}-m)=1$. If $m$ is odd, then $\nu_{2}(m)=0$ and $\nu_{2}(m-1)=1$ iff $m\equiv 3\pmod{4}$. Thus among square-free $m$, the condition is $m\not\equiv 1\pmod{4}$. The density follows from Theorem 3.6 with $p=2$.
(2) Here $p\in\{2,3\}$. The $p=2$ condition is as in (1). For $p=3$ note $m^{3}-m=m(m-1)(m+1)$. If $3\mid m$ and $m$ is square-free, then $\nu_{3}(m^{3}-m)=1$. If $3\nmid m$, then $\nu_{3}(m^{3}-m)=1$ unless $m\equiv\pm 1\pmod{9}$, in which case $\nu_{3}\geq 2$. Combining gives the claimed congruences, and the density follows from Theorem 3.6. ∎
Remark 3.8 (Cubic case and field‑monogeneity).
For cube‑free $m$ and $p=3$, we have $\mathbb{Z}[\sqrt[3]{m}]=\mathcal{O}_{K}$ iff $m\not\equiv\pm 1\pmod{9}$. In the exceptional cases $m\equiv\pm 1\pmod{9}$, the field is nevertheless monogenic (though not via $\alpha$): one can take
$\theta=\frac{1\pm\alpha+\alpha^{2}}{3}.$
It is classical that $\theta$ is an algebraic integer in these residue classes and that $\mathcal{O}_{K}=\mathbb{Z}[\theta]$.
3.3. Consequences and applications of the density formula
Write $\delta_{n}=\frac{6}{\pi^{2}}\prod_{p\mid n}\frac{p}{p+1}$, we obtain the global counting with error term as follows.
Corollary 3.9.
Fix $n\geq 2$. Let
$\mathcal{M}_{n}(X)=\#\{\,|m|\leq X:\ x^{n}-m\text{ irreducible and }\mathcal{O}_{K}=\mathbb{Z}[\alpha]\,\}.$
Then
$\mathcal{M}_{n}(X)=2\delta_{n}\,X+O_{n}\!\left(X^{1/2}\right).$
If one restricts to $1\leq m\leq X$ the main term is $\delta_{n}X$ with the same error term.
Proof.
By Theorem 3.6 the relevant set has natural density $\delta_{n}$. The $O(X^{1/2})$ error follows from the proof of Lemma 3.1 (square-free counts in classes) combined with inclusion–exclusion over the finitely many primes $p\mid n$. ∎
The following result describes the arithmetic progressions together with their explicit local factors and depths.
Corollary 3.10.
Let $q\geq 1$ and $a\in\mathbb{Z}$ with $\gcd(a,q)=1$. Write $v_{p}(q)$ for the pp‑adic valuation. Then, among integers $m\equiv a\bmod q$ for which $x^{n}-m$ is irreducible, the set of $m$ with $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ has natural density
$\frac{1}{q}\cdot\frac{1}{\zeta(2)}\cdot\prod_{\ell\mid q}\frac{1}{1-\ell^{-2}}\ \times\ \prod_{p\mid n}\lambda_{p}(q,a),$
where the local factor at $p\mid n$ is
$\lambda_{p}(q,a)=\begin{cases}\displaystyle\frac{p}{p+1},&v_{p}(q)=0,\\[6.0pt] \displaystyle\frac{p-1}{p},&v_{p}(q)=1,\\[6.0pt] 1,&v_{p}(q)\geq 2\ \text{and}\ a\not\in\mathcal{E}_{p}\ \text{mod }p^{2},\\[4.0pt] 0,&v_{p}(q)\geq 2\ \text{and}\ a\in\mathcal{E}_{p}\ \text{mod }p^{2},\end{cases}$
with $\mathcal{E}_{p}=\{x\bmod p^{2}:\ x^{p}\equiv x\}$. For $p=2$ interpret this modulo $4$ (the unique bad unit class is $1\bmod 4$).
Proof.
By Lemma 3.1, among $m\equiv a\bmod q$ the square‑free density is $\frac{1}{q\zeta(2)}\prod_{\ell\mid q}(1-\ell^{-2})^{-1}$. Fix $p\mid n$. [^1] Uniformity among lifts from a fixed residue modulo $p$ to residues modulo $p^{2}$ follows by applying Lemma 3.1 to each class modulo $p^{2}$ and summing, using the Chinese Remainder Theorem to separate the $p$-part from the rest of the modulus.
If $v_{p}(q)=0$ there is no local information at $p$, so the survival probability is the global one $p/(p+1)$ (Proposition 3.4).
If $v_{p}(q)=1$, the congruence fixes $m\bmod p$ with $p\nmid a$. Among the $p$ lifts modulo $p^{2}$ above $a\bmod p$, exactly one is Teichmüller (bad), and square‑free integers are uniformly distributed across these lifts (by Lemma 3.1 with modulus $p^{2}$). Hence survival is (p1)/p(p-1)/p (Proposition 3.17).
If $v_{p}(q)\geq 2$, then $m\equiv a\bmod p^{2}$ is fixed. If $a\bmod p^{2}\in\mathcal{E}_{p}$, every such $m$ fails at $p$, otherwise every such $m$ survives at $p$. (For $p=2$ it suffices to work modulo $4$ for the valuation $\nu_{2}(m^{2}-m)$.)
Independence over $p$ (by the Chinese Remainder Theorem) yields the product of the local factors. ∎
Example 3.11 (Quartic case in a fixed progression).
Let $n=4$, $q=4$, $a\equiv 3\bmod 4$. Here $v_{2}(q)=2$ and $a\not\equiv 1\bmod 4$, so $\lambda_{2}(q,a)=1$. Therefore
$\delta\big(\{m\equiv 3\bmod 4:\ \mathcal{O}_{K}=\mathbb{Z}[\sqrt[4]{m}]\}\big)=\frac{1}{4}\cdot\frac{1}{\zeta(2)}\cdot\frac{1}{1-2^{-2}}=\frac{2}{\pi^{2}}.$
This reflects that on the class $3\bmod 4$ the $p=2$ constraint is automatically satisfied, only square‑freeness remains.
Remark 3.12 (Canonical modulus for refinement).
For questions beyond $\alpha$-monogeneity (e.g. tracking the full $p$-denominator profile of an integral basis), the canonical modulus is $M(n)=\prod_{p^{e}\parallel n}p^{e+1}=n\cdot\operatorname{rad}(n)$: the local $p$-shape depends only on $a\bmod p^{e+1}$ and this modulus is minimal. Our $\alpha$-monogeneity constraints at $p\mid n$ depend only on $a\bmod p^{2}$, but working modulo $M(n)$ aligns directly with the periodicity tables of the full shape, see [NGU25].
The following result exhibits arithmetic progressions of square-free $m$ with $100\%$ $\alpha$-monogeneity.
Corollary 3.13.
Define
$M=\begin{cases}\displaystyle\prod_{\begin{subarray}{c}p\mid n\\ p\ \mathrm{odd}\end{subarray}}p^{2},&2\nmid n,\\[6.0pt] \displaystyle 4\prod_{\begin{subarray}{c}p\mid n\\ p\ \mathrm{odd}\end{subarray}}p^{2},&2\mid n.\end{cases}$
There exist residue classes $a\bmod M$ with $\gcd(a,M)=1$ such that every square-free $m\equiv a\bmod M$ satisfies $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$. Consequently, for each such $a$ the set $\{m\equiv a\bmod M:\ \mathcal{O}_{K}=\mathbb{Z}[\alpha]\}$ has density
$\frac{1}{M}\cdot\frac{1}{\zeta(2)}\cdot\prod_{\ell\mid M}\frac{1}{1-\ell^{-2}}.$
Proof.
Choose $a$ avoiding the finitely many bad classes modulo $M$: if $2\mid n$, avoid $1\bmod 4$ (and $\gcd(a,4)=1$ forces $a$ odd), and for each odd $p\mid n$ avoid the (p1)(p-1) Teichmüller classes modulo $p^{2}$. Then for any square-free $m\equiv a\bmod M$ one has $\nu_{p}(m^{p}-m)=1$ for all $p\mid n$. The density among $m\equiv a\bmod M$ is then given by Lemma 3.1. ∎
We obtain the following theorem, which gives the asymptotic count of $\alpha$-monogenic pure fields ordered by discriminant.
Theorem 3.14.
Let $N_{n}(Y)$ be the number of fields $K=\mathbb{Q}(\sqrt[n]{m})$ with $|\,\Delta_{K}\,|\leq Y$, $x^{n}-m$ irreducible, and $\mathcal{O}_{K}=\mathbb{Z}[\sqrt[n]{m}]$. Then, with $X=(Y/n^{n})^{1/(n-1)}$,
$N_{n}(Y)=\begin{cases}\ \delta_{n}\,X\ +\ O_{n}\!\big(X^{1/2}\big),&n\ \text{odd},\\[4.0pt] \ 2\delta_{n}\,X\ +\ O_{n}\!\big(X^{1/2}\big),&n\ \text{even}.\end{cases}$
Proof.
When $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$, the field discriminant equals the polynomial discriminant by (2.1): $\Delta_{K}=\mathrm{disc}(x^{n}-m)=\pm n^{n}m^{\,n-1}$. Thus $|\,\Delta_{K}\,|\leq Y$ is equivalent to $|m|\leq X$. For $n$ odd, the fields with parameters $m$ and m-m coincide (Indeed, when $n$ is odd, replacing $\alpha$ by $\alpha$ leaves $\mathbb{Q}(\alpha)$ unchanged, when $n$ is even, $\mathbb{Q}(\sqrt[n]{m})\neq\mathbb{Q}(\sqrt[n]{-m})$.) Hence counting positive $m$ gives distinct fields and yields the factor $\delta_{n}$, for $n$ even, $m\mapsto-m$ gives distinct fields, hence the factor $2\delta_{n}$. Apply Corollary 3.9. ∎
Immediate from $\delta_{n}=(6/\pi^{2})\prod_{p\mid n}p/(p+1)$ and the fact that adjoining a new prime factor multiplies by $p/(p+1)<1$ we obtain the following.
Example 3.15.
If $n\mid n^{\prime}$ then $\delta_{n^{\prime}}\leq\delta_{n}$. Explicitly,
$\delta_{2}=\delta_{4}=\frac{4}{\pi^{2}}\approx 0.4053,\qquad\delta_{3}=\frac{9}{2\pi^{2}}\approx 0.4559,\qquad\delta_{6}=\frac{3}{\pi^{2}}\approx 0.3040,\qquad\delta_{5}=\frac{5}{\pi^{2}}\approx 0.5066.$
The following result describes the failure distribution and independence phenomena.
Proposition 3.16.
Let $P=\{p:\ p\mid n\}$. Among square-free $m$, the events $\{m\in\mathcal{E}_{p}\}$ for $p\in P$ are independent. Hence, for any subset $S\subseteq P$,
[…]
$=\frac{6}{\pi^{2}}\Big(\prod_{p\in S}\frac{1}{p+1}\Big)\Big(\prod_{p\in P\setminus S}\frac{p}{p+1}\Big).$
In particular, the expected number of failing primes (among those dividing $n$) equals $\sum_{p\mid n}\frac{1}{p+1}$.
Proof.
By (3.1) with CRT, for any finite $P$ we have
$\delta\!\left(S\cap\bigcap_{p\in P}\mathcal{E}_{p}\right)=\frac{1}{\zeta(2)}\prod_{p\in P}\frac{p-1}{p^{2}-1}=\left(\frac{1}{\zeta(2)}\right)\prod_{p\in P}\left(\frac{1}{p+1}\cdot\frac{p}{p-1}\cdot\frac{p-1}{p}\right)=\delta(S)\prod_{p\in P}\frac{1}{p+1},$
so the events are independent within $S$ (and intersecting with the density‑1 irreducible set does not change the density). The stated formula and the expected value follow. ∎
Proposition 3.17 (Conditional local laws at a fixed prime).
Let $p\mid n$ and fix a residue class $a\!\!\pmod{p}$. Among square-free $m\equiv a\bmod p$, the survival probability at $p$ equals:
$\begin{cases}1,&a\equiv 0\ (\bmod p),\\[4.0pt] \frac{p-1}{p},&a\not\equiv 0\ (\bmod p).\end{cases}$
Consequently, conditioning on $m\equiv 0\pmod{p}$ neutralizes the $p$-constraint, while conditioning on any $a\not\equiv 0$ replaces the global factor $p/(p+1)$ by (p1)/p(p-1)/p.
Proof.
If $a\equiv 0$, then for square-free $m$ one has $v_{p}(m)=1$ and so $v_{p}(m^{p}-m)=1$ automatically. If $a\not\equiv 0$, then among the $p$ lifts modulo $p^{2}$ lying above $a$ exactly one is Teichmüller (bad), and by Lemma 3.1 each such lift carries the same square-free density, hence the survival fraction is (p1)/p(p-1)/p. ∎
Remark 3.18 (From $p$-level to $p^{2}$-level conditioning).
Proposition 3.17 conditions only on $m\bmod p$, giving survival probability (p1)/p(p-1)/p when $a\not\equiv 0\bmod p$. If one instead fixes $m\bmod p^{2}$, the survival probability becomes either $1$ (if $a\bmod p^{2}\notin\mathcal{E}_{p}$) or 0 (if $a\bmod p^{2}\in\mathcal{E}_{p}$). This is the mechanism behind Corollary 3.13: choosing a modulus $M$ divisible by $p^{2}$ for every $p\mid n$ (and $4$ for $p=2$) and selecting a good class modulo each $p^{2}$ forces the pp–constraint to hold identically on the progression.